# Structure and Independence in Hyperbolic Uniform Disk Graphs

### v1

@@ 2 Preliminaries
The Voronoi cells are convex. Thus, drawing each edge $\{u,v\}$ of the Delaunay complex $\mathcal{D}(S)$ by choosing an internal point $p$ on the Voronoi edge dual to $\{u,v\}$ and then connecting $u$ via $p$ to $v$ using the two line segments $up$ and $pv$ yields a planar drawing of $\mathcal{D}(S)$; see Figure 3(c). It has the property that each edge of $\mathcal{D}(S)$ intersects its dual Voronoi edge in exactly one point and it intersects no other Voronoi edges.
@@ 3 Balanced separators in HUDGs
Let $G$ be a hyperbolic uniform disk graph with $n$ vertices and radius $r$. We show that $G$ has a balanced separator that can be covered by $\mathcal{O}((1+1/r)\log n)$ cliques; see Theorem 1. The overall argument is as follows. We find a double wedge bounded by two lines $\ell_{1}$ and $\ell_{2}$ that contains no vertex in its interior and separates the other vertices of $G$ in a balanced fashion. In Figure 4(a), the double wedge with apex $p$ is shaded gray and contains no vertices and the regions above and below the wedge each contain a constant fraction of the vertices.
Let $m$ be the angular bisector of the wedge. As separator, we use the set of all vertices that have distance at most $r$ from $m$. The set of points with distance exactly $r$ from $m$ forms two curves that are called hypercycles with axis $m$. Thus, vertices belong to the separator if they lie on or between the two hypercycles. The hypercycles are shown in Figure 4(a) and the region where separator vertices can lie is shaded blue. In the Poincaré disk, hypercycles are arcs of Euclidean circles that meet the boundary of the disk at the same ideal points as their axis but at a non-right angle. Observe that any vertex from above the top hypercycle has distance greater than $2r$ to any vertex below the bottom hypercycle. Thus, the vertices in the blue region indeed form a balanced separator.
It remains to show that the graph induced by vertices in the blue region can be covered with few cliques. For this, we cover the blue region with boxes as shown in Figure 4(b). We show that each of these boxes has diameter at most $2r$, implying that the vertices within one box induce a clique. Moreover, we show that we only need $\mathcal{O}((1+1/r)\log n)$ such boxes. For the latter, we need a lower bound on the opening angle of the empty double wedge. Observe that a larger opening angle $\varphi$ in Figure 4(c) shrinks the blue region, which reduces the number of boxes required to cover it.
In Section 3.1, we give a bound on the number of cliques required to cover the separator, parameterized by the opening angle of the empty double wedge. Afterwards, in Section 3.2 we show that there is a wedge with not too small opening angle that yields a balanced separator. The proof is constructive, i.e., given the graph with vertex positions, we can efficiently find the wedge and thereby a balanced separator.
Figure 4: Illustration of our separator using the Poincaré disk model. (4(a)) The axis $m$ is chosen such that it separates the vertices in a balanced fashion. The double wedge between $\ell_{1}$ and $\ell_{2}$ contains no vertices. The separator consists of the blue set of all vertices of distance at most $r$ to the axis $m$. (4(b)) We cover the separator with boxes that have diameter $2r$ and thus form cliques. (4(c)) If the opening angle $\varphi$ is sufficiently large, then $x$ is sufficiently small and we only need few boxes to cover the whole separator.
@@ 3.1 Covering the separator with cliques
We start by defining the boxes between a hypercycle and its axis $m$ as illustrated in Figure 4(b). We place points $b_{0},\dots,b_{k}$ along $m$, such that $b_{0}=p$ and they are evenly spaced at distance $\tanh(r)$. Through each point $b_{i}$ we draw a line perpendicular to $m$ and let $a_{i}$ denote its intersection with the upper hypercycle, see Figure 4(b). We call the region bounded by the line segments $a_{i-1}b_{i-1}$, $b_{i-1}b_{i}$, $b_{i}a_{i}$, and the part of the hypercycle between $a_{i}$ and $a_{i-1}$ a box.
{lemma}
@@
To calculate the distance from $a_{0}$ to $a_{1}$, we can use that the polygon $a_{0},b_{0},b_{1},a_{1}$ is a Saccheri quadrilateral with base $b_{0}b_{1}$ of length $\tanh r$, legs $b_{0}a_{0}$ and $b_{1}a_{1}$ of length $r$, and summit $a_{0}a_{1}$. The length of the summit is given by the following formula [Euclid_Non_Euclid_Geomet-Green93, Theorem 10.8]:
@@
since $\sinh r/\cosh r=\tanh r>\tanh 1>\sinh(1/2)$ as $\tanh r$ is monotone increasing. Consequently, $|a_{0}a_{1}|<2r$. Thus, all distances between vertices of a box are at most $2r$, which concludes the proof. ∎
From this lemma, it follows that the vertices inside any box induce a clique. Next, we aim to give an upper bound on the number of boxes we need to cover the separator. For this, let $k$ be the smallest number such that $a_{k}$ lies in the wedge formed by $\ell_{1}$ and $\ell_{2}$. In Figure 4(b), $k=3$ as $a_{3}$ lies in the gray wedge. Clearly, we can cover the whole separator (blue in Figure 4), with $\mathcal{O}(k)$ boxes on both sides of the line and thus with $\mathcal{O}(k)$ cliques. The following lemma gives a bound on $k$ depending on the opening angle of the wedge.
{lemma}
@@
Proof.
Consider the triangle in Figure 4(c), where $p=b_{0}$ is the apex of the wedge, $d$ is the intersection of the wedge boundary $\ell_{1}$ with the hypercycle, and $c$ is the point on the axis with distance $r$ from $d$. Recall that the distance between $b_{i-1}$ and $b_{i}$ is $\tanh r<1$. Thus, the distance between $b_{0}$ and $b_{k}$ equals the distance $x$ between $p$ and $c$ up to an additive constant. It remains to give an upper bound on $x$.
@@
Using that $\tanh r<1$ and $\tan\varphi>\varphi$, we obtain $\sinh x<\frac{1}{\varphi}$. This yields the claim. ∎

### v2

@@ 2 Preliminaries
The Voronoi cells are convex. Thus, drawing each edge $\{u,v\}$ of the Delaunay complex $\mathcal{D}(S)$ by choosing an internal point $p$ on the Voronoi edge dual to $\{u,v\}$ and then connecting $u$ via $p$ to $v$ using the two line segments $up$ and $pv$ yields a planar drawing of $\mathcal{D}(S)$; see Figure 3. It has the property that each edge of $\mathcal{D}(S)$ intersects its dual Voronoi edge in exactly one point and it intersects no other Voronoi edges.
3 Overview of our techniques
3.1 Balanced separators in HUDGs
Let $G$ be a hyperbolic uniform disk graph with $n$ vertices and radius $r$. We show that $G$ has a balanced separator that can be covered by $\mathcal{O}((1+1/r)\log n)$ cliques; see Theorem 1.1. The overall argument is as follows. We find a double wedge bounded by two lines $\ell_{1}$ and $\ell_{2}$ that contains no vertex in its interior and separates the other vertices of $G$ in a balanced fashion, which is ensured by selecting the center of the double wedge to be the hyperbolic centerpoint [28]. In Figure 4, the double wedge with apex $p$ is shaded gray and contains no vertices and the regions above and below the wedge each contain a constant fraction of the vertices.
Let $m$ be the angular bisector of the wedge. As separator, we use the set of all vertices that have distance at most $r$ from $m$. The set of points with distance exactly $r$ from $m$ forms two curves that are called hypercycles with axis $m$. Thus, vertices belong to the separator if they lie on or between the two hypercycles. The hypercycles are shown in Figure 4 and the region where separator vertices can lie is shaded blue. In the Poincaré disk, hypercycles are arcs of Euclidean circles that meet the boundary of the disk at the same ideal points as their axis but at a non-right angle. Observe that any vertex from above the top hypercycle has distance greater than $2r$ to any vertex below the bottom hypercycle. Thus, the vertices in the blue region indeed form a balanced separator.
It remains to show that the graph induced by vertices in the blue region can be covered with few cliques. For this, we cover the blue region with boxes as shown in Figure 4. We show that each of these boxes has diameter at most $2r$, implying that the vertices within one box induce a clique. Moreover, we show that we only need $\mathcal{O}((1+1/r)\log n)$ such boxes. For the latter, we need a lower bound on the opening angle of the empty double wedge. Observe that a larger opening angle, $\varphi$, in Figure 4, shrinks the blue region, which reduces the number of boxes required to cover it.
In Section 5.1, we give a bound on the number of cliques required to cover the separator, parameterized by the opening angle of the empty double wedge. In Section 5.2 we show that there is a wedge with not too small opening angle that yields a balanced separator. The proof is constructive, i.e., given the graph with vertex positions, we can efficiently find the wedge and thereby the balanced separator of Theorem 1.1.
Figure 4: Illustration of our separator using the Poincaré disk model. (4) The axis $m$ is chosen such that it separates the vertices in a balanced fashion. The double wedge between $\ell_{1}$ and $\ell_{2}$ contains no vertices. The separator consists of the blue set of all vertices of distance at most $r$ to the axis $m$. (4) We cover the separator with boxes that have diameter $2r$ and thus form cliques. (4) If the opening angle $\varphi$ is sufficiently large, then $x$ is sufficiently small and we only need few boxes to cover the whole separator.
3.2 Outerplanarity of hyperbolic Delaunay complexes
Let $S$ be a set of $n$ sites in $\mathbb{H}^{2}$ with pairwise distance at least $2r$. Note that $S$ interpreted as a hyperbolic uniform disk graph with radius $r$ forms an independent set. To prove Theorem 1.4, giving an upper bound on the outerplanarity of the Delaunay complex $\mathcal{D}(S)$, we use two different types of arguments. The first argument is based on the observation that sufficiently large radius implies that any inner vertex of $\mathcal{D}(S)$ requires many other sites around it to shield it from being an outer vertex. This is similar to the observation in the introduction (recall Figure 1) that for high radius $r$, large stars can be realized. The argument then goes roughly as follows: If all inner vertices have degree more than $6$, then we need a linear fraction of vertices on the outer face to account for the fact that the average degree in planar graphs is bounded by $6$. This already implies that iteratively deleting the vertices on the outer face takes at most $\log n$ steps as each step deletes a constant fraction of the vertices. A larger radius $r$ leads to higher degrees of inner vertices, which in turn yields stronger bounds for the outerplanarity.
This type of argument can only work if the radius $r$ is sufficiently large. For smaller radii, the second type of argument considers layers of bounded Voronoi cells (which correspond to inner vertices of $\mathcal{D}(S)$) around one fixed center cell. As the union of these layers is bounded by a polygon with a linear number of vertices, its area is linear. In contrast to that, we will see that the area of the layers grows exponentially with the layer (with the base depending on $r$), showing that there cannot be too many layers.
In the following, we give a sketch of how the second argument based on area works for small radii. Although this is the argument of choice for radii that are small constants or even decreasing with $n$, we first give a simple area-based argument for why $r\geq\log n$ implies that the Delaunay complex has no inner vertices. This gives a good intuition how the area behaves in the hyperbolic plane and how this can help us in proving outerplanarity, even though our argument for small radii is more involved. Let $s\in S$ be a site. If $s$ is an inner vertex of $\mathcal{D}(S)$, then its Voronoi cell is bounded and its boundary is a polygon of at most $n-1$ vertices. Thus, the area of this Voronoi cell is less than $(n-3)\pi$. Simultaneously, the disk of radius $r$ around $s$ is included in the Voronoi cell of $s$ as all sites have pairwise distance at least $2r$. The area of this disk is $4\pi\sinh^{2}(\frac{r}{2})$ which is larger than $e^{r}$ for sufficiently large $r$. From this it follows that bounded cells can only exist when $r<\log n$.
To make an argument that works for smaller $r$, let $s\in S$ be an arbitrary but fixed vertex of the Delaunay complex $\mathcal{D}(S)$. We partition the vertices of $\mathcal{D}(S)$ into layers by hop distance from $s$ in $\mathcal{D}(S)$, i.e., $V_{\ell}$ is the set of vertices with distance $\ell$ from $s$. Let $L$ be the largest integer such that for all $\ell\leq L$ the layer $V_{\ell}$ contains only inner vertices. Note that our goal is to prove an upper bound on $L$, as this bounds the distance of $s$ to the outer face.
The details of this area-based argument for small radii can be found in Section 6.2. Together with the degree-based bound for large radii, see Section 6.1 for details, we obtain Theorem 1.4.
3.3 Exact algorithm
Our exact algorithm stated in Theorem 1.5 works roughly as follows; see Section 7 for details. Instead of computing the independent set directly, we define some additional structure for a given independent set. The algorithm then finds this additional structure such that the corresponding independent set is maximized. The additional structure is illustrated for two example independent sets in Figure 6. In the following, we explain this from top to bottom.
Figure 6: Illustration of our exact algorithm in Theorem 1.5.
Let $G$ be a hyperbolic uniform disk graph of radius $r$ together with an independent set $S$ (top row in Figure 6). As there are no edges between vertices in $S$, they have pairwise distance at least $2r$. Thus, the Delaunay complex $\mathcal{D}(S)$ (second row in Figure 6) has low outerplanarity due to Theorem 1.4. Low outerplanarity implies small treewidth, which implies small branchwidth [43]. Moreover, as $\mathcal{D}(S)$ is a planar graph, we can wish for a so-called sphere cut decomposition of low width, which is a branch decomposition with some additional properties [44, 17, 37]. In a nutshell, we get a hierarchy of nooses, where each noose is a closed curve that intersects the graph only at vertices and functions as a vertex separator, splitting interior edges from exterior ones. The hierarchy is a binary tree in the sense that every noose containing more than one edge in its interior has two child nooses partitioning these edges further. Row three in Figure 6 shows three nooses of the Delaunay complex $\mathcal{D}(S)$. The solid blue noose is the parent of the dashed red and green noose.
The width of a sphere cut decomposition is the maximum number of vertices on a noose. As $\mathcal{D}(S)$ has low outerplanarity, we have a low-width sphere cut decomposition, i.e., each noose goes through only few vertices. As a last step, we normalize the nooses by specifying their exact geometry, keeping their combinatorial structure intact; see the last row of Figure 6. At its core, the normalized nooses are polygons that alternate between vertices in $S$ and Voronoi vertices, with some special treatment for the outer face of $\mathcal{D}(S)$.
Observe that the two examples (left and right column) illustrate different independent sets $S$, but share the green normalized noose. This is not a coincidence, but actually happens a lot. In fact, despite the potentially exponential number of independent sets, we can bound the total number of normalized nooses that we can obtain in this way by $n^{\mathcal{O}(1+\frac{\log n}{r})}$. This follows from the fact that each noose visits only $\mathcal{O}(1+\frac{\log n}{r})$ vertices due to Theorem 1.4. Thus, the normalized nooses are essentially polygons with $\mathcal{O}(1+\frac{\log n}{r})$ corners and the set of potential corners is polynomial, as each corner is a vertex of $G$ or a Voronoi vertex (which is defined by three vertices of $G$).
Our independent set algorithm works as follows. We start by listing all candidates for normalized nooses. We then use a dynamic program to combine nooses into a hierarchy of polygons. We make sure that the resulting hierarchy is valid in the sense that it corresponds to a sphere cut decomposition. Additionally, we enforce that it is well-spaced, which lets us show that all vertices of $G$ visited by the polygons in the hierarchy form an independent set. Then, finding a maximum independent set is equivalent to maximizing the number of vertices visited by polygons in this polygon hierarchy, which can be done with the dynamic program.
3.4 Approximation algorithm
The approximation algorithm of Theorem 1.6 (fully described in Section 8) works similarly to Lipton and Tarjan’s version for planar graphs [33]: we repeatedly apply a separator until we get small patches and then solve each patch individually using the exact algorithm of Theorem 1.5. Vertices in the separators get ignored, so we get an $(1-\varepsilon)$-approximation when these make up only an $\varepsilon$ fraction of the maximum independent set. We guarantee this using a lower bound of $\Omega(n/\ell)$ on the size of a maximum independent set, which follows from a proof that hyperbolic uniform disk graphs of ply $\ell$ always have a vertex of degree at most $14\ell-$ 1, making them $(4\ell-1)$-degenerate.
4 Grids, and stars in hyperbolic uniform disk graphs
In this section we revisit a claim from the introduction and show that hyperbolic uniform disk graphs with radius $1/n^{3}$ and radius $\log n$ are incomparable. The former class contains grids but no large stars while the latter contains all stars but no large grids.
Theorem 4.1.
Let $\Gamma_{k}$ denote the $k\times k$ grid graph on $n=k^{2}$ vertices, and let $S_{n}$ denote the star on $n$ vertices. Then we have the following:
$\Gamma_{k}\in\mathrm{HUDG}(1/n^{3})$ for all $k$,
$S_{n}\not\in\mathrm{HUDG}(1/n^{3})$ for $n\geq 8$,
$\Gamma_{k}\not\in\mathrm{HUDG}(\log n)$ for $k\geq 5$,
$S_{n}\in\mathrm{HUDG}(\log n)$ for all $n$.
Proof 4.2.
(i) Consider a mapping from grid vertices to coordinates in the Poincaré half-plane model, where the vertex $v=(i,j)$ is mapped to the coordinate $p_{v}=(2i/n^{3},1+2j/n^{3})\in\mathbb{R}\times\mathbb{R}_{>0}$ for $(i,j)\in[k]\times[k]$. For a detailed introduction to the upper half-plane model see [42, p. 128]. The distance between two points $p_{u}=(x_{u},y_{u})$ and $p_{v}=(x_{v},y_{v})$ in the upper half-plane can be expressed by $d_{\mathbb{H}^{2}}(p_{u},p_{v})=\mathrm{arcosh}(1+\frac{\left\lVert p_{u}-p_{v}\right\rVert^{2}}{2y_{u}y_{v}})$.
For two grid vertices $u=(i,j)$, $v=(i^{\prime},j^{\prime})$, we first consider the terms $\left\lVert p_{u}-p_{v}\right\rVert^{2}$ and $2y_{u}y_{v}$ appearing in the distance function. We start by giving a lower and upper bound for $2y_{u}y_{v}$. Observe that $y_{u},y_{v}\geq 1$ and thus $2y_{u}y_{v}\geq 2$. Conversely, as the $j,j^{\prime}\leq\sqrt{n}$, we have $y_{u},y_{v}\leq 1+2\sqrt{n}/n^{3}$. As we can assume $n\geq 4$, we get $y_{u},y_{v}\leq 1+\frac{1}{16}$ and thus $32y_{u}y_{v}\leq 2(1+\frac{1}{16})^{2}<$ 3. Thus, $32\leq 2y_{u}y_{v}<$ 3. For the term $\left\lVert p_{u}-p_{v}\right\rVert^{2}$ we get
$\left\lVert p_{u}-p_{v}\right\rVert^{2}=\frac{4(i-i^{\prime})^{2}}{n^{6}}+\frac{4(j-j^{\prime})^{2}}{n^{6}}=\frac{4}{n^{6}}\left((i-i^{\prime})^{2}+(j-j^{\prime})^{2}\right).$
In case $u$ and $v$ are adjacent, one of $(i-i^{\prime})^{2}$ and $(j-j^{\prime})^{2}$ is $1$ and the other is 00 and thus we get $\left\lVert p_{u}-p_{v}\right\rVert^{2}=4/n^{6}$. In case $u$ and $v$ are not adjacent, we get $(i-i^{\prime})^{2}+(j-j^{\prime})^{2}\geq 2$ and thus $\left\lVert p_{u}-p_{v}\right\rVert^{2}\geq 8/n^{6}$. Together with the bound $32\leq 2y_{u}y_{v}<$ 3, we obtain
$\frac{\left\lVert p_{u}-p_{v}\right\rVert^{2}}{2y_{u}y_{v}}\leq\frac{4}{2n^{6}}=\frac{2}{n^{6}}\text{ if }\{u,v\}\in E\quad\text{ and }\quad\frac{\left\lVert p_{u}-p_{v}\right\rVert^{2}}{2y_{u}y_{v}}>\frac{8}{3n^{6}}\text{ if }\{u,v\}\notin E.$
As $\mathrm{arcosh}$ is strictly increasing (and $2<8/3$), this already shows that there exists a disk radius $r$ such that we get a uniform disk representation of $\Gamma_{k}$. It remains to show that specifically $r=1/n^{3}$ works.
For this, we use that the series expansion of $\mathrm{arcosh}(1+x)$ gives
$\sqrt{2x}-\frac{x^{3/2}}{6\sqrt{2}}\leq\mathrm{arcosh}(1+x)\leq\sqrt{2x}.$
If $u$ and $v$ are adjacent, this yields $d_{\mathbb{H}^{2}}(p_{u},p_{v})\leq\mathrm{arcosh}(1+2/n^{6})\leq 2/n^{3}$. Thus, the two disks with radius $1/n^{3}$ intersect. If $u$ and $v$ are not adjacent, we obtain
$d_{\mathbb{H}^{2}}(p_{u},p_{v})>\mathrm{arcosh}(1+\frac{8}{3n^{6}})\geq\sqrt{2\frac{8}{3n^{6}}}-\frac{\left(\frac{8}{3n^{6}}\right)^{3/2}}{6\sqrt{2}}>\frac{2.3}{n^{3}}-\frac{0.6}{n^{9}}.$
For $n\geq 4$, this is clearly greater than $2/n^{3}$ and thus the two disks of radius $1/n^{3}$ do not intersect. Thus, the above defined coordinates yield an intersection representation of $\Gamma_{k}$ with disk of uniform radius $r=1/n^{3}$.
(ii) Suppose for contradiction that $S_{n}$ has a uniform disk representation with radius $r=1/n^{3}$ for $n\geq 8$. Let $c$ be the center vertex of $S_{n}$. As $S_{n}$ has at least $7$ leaves, there are at least two leaves $u$ and $v$ such that the angle between the line segments $cu$ and $cv$ is at most $72\pi/$ 7. We will show, that the distance $x$ between $u$ and $v$ is at most $2r$; a contradiction. As $c$ is connected to $u$ and $v$, both have distance at most $2r$ from $c$. We first make an argument that the distance between $u$ and $v$ is maximized if either both have distance $2r$ from $c$ or if one has distance $2r$ and the other has the same position as $c$. Assume we fix the position of $v$ and move $u$ along the line through $u$ and $c$. Then the distance of $u$ to $v$ is unimodal with a unique minimum and moving $u$ in either direction from the minimum only increases the distance. Thus, the distance between $u$ and $v$ is maximized by moving $u$ such that it has maximum distance from $c$ or is equal to $c$.
If $u$ has the same position as $c$, then $u$ and $v$ would be connected. Thus, it remains to consider the case where $u$ and $v$ both have distance $2r$ from $c$ and thus form an isosceles triangle with side-length $2r$ and angle $\alpha\leq 2\pi/7$; see Figure 7. This triangle consists of two right triangles for which we can apply the sine formula for hyperbolic triangles to get
$\sin\frac{\alpha}{2}=\frac{\sinh\frac{x}{2}}{\sinh 2r}\quad\Leftrightarrow\quad\sinh\frac{x}{2}=\sin\frac{\alpha}{2}\cdot\sinh 2r.$
Note that $u$ and $v$ are adjacent if and only if $x\leq 2r$, from which we can derive the following condition.
$x\leq 2r\quad\Leftrightarrow\quad\sinh\frac{x}{2}\leq\sinh r\quad\Leftrightarrow\quad\sin\frac{\alpha}{2}\cdot\sinh 2r\leq\sinh r.$
As $\alpha\leq 2\pi/7$, we get $\sin\frac{\alpha}{2}\leq 0.44$. Moreover, $\sinh r/\sinh 2r$ is a decreasing function for $r>0$ and it holds that $\sinh r/\sinh 2r>0.49$ for $r\leq 1/8^{3}$. Thus, the above inequality $\sin\frac{\alpha}{2}\cdot\sinh 2r\leq\sinh r$ holds and hence $x\leq 2r$, which means that $u$ and $v$ are connected; a contradiction.
Figure 7: Visualizations for the proof of Theorem 4.1: (a) drawing of center $c$ and leaves $v$, $w$ together with the relevant right triangle in an enlarged detail view of the Poincaré disk model, relevant in part (ii); (b) subgraph $H$ of the $5\times 5$ grid with highlighted degree $3$ vertices, relevant in part (iii).
(iii) We define a subgraph $H$ as follows, see also Figure 7. Consider the standard Euclidean drawing of the $5\times 5$ grid and let $C$ denote the cycle of length 16 given by the set of edges on the outer face, and let $v$ be the vertex in the center of the grid. Consider $C$ and the $2$-hop paths connecting $v$ to the middle vertices of the top, left, and right sides of the square of $C$. These paths together with $C$ form a subdivision of a $4$-clique; let $H$ be this subgraph of $\Gamma_{5}$.
Consider now a realization of $\Gamma_{k}$ for $k\geq 5$ with hyperbolic disks of radius $\log n$. We find $H$ as an induced subgraph of $\Gamma_{k}$ (as $\Gamma_{5}$ is a subgraph of $\Gamma_{k}$) and the corresponding disks in the realization of $\Gamma_{k}$ realize $H$. Consider the drawing of $H$ obtained by connecting the centers of adjacent disks in $H$ with straight line segments. This drawing is planar for the following reason. Assume two edges $\{a,b\}$ and $\{c,d\}$ intersect in a point $x$ and assume without loss of generality that $a$ is the vertex among $\{a,b,c,d\}$ that is closest to $x$. Then it follows from the triangular inequality that $a$ must have an edge to $c$ and $d$, a contradiction to the fact that $H$ has no triangles.
Thus, we obtain a plane straight-line drawing of a subdivided $4$-clique and one of the vertices of $H$ (say, $w$) of degree $3$ cannot be on the outer face. Notice that the outer face is bounded by a cycle $C_{w}$ that is disjoint from the neighborhood of $w$. Thus, $C_{w}$ forms a polygon and the disk of $w$ must be fully contained in this polygon. Since $C_{w}$ has at most $16$ vertices, the area of its polygon is at most that of $14$ hyperbolic triangles, i.e., at most $14\pi$, which is less than the area of a disk of radius $\log n$ if $n\geq 25$; a contradiction.
(iv) We may assume $n\geq 3$ and realize $S_{n}$ with radius $r=\log n$, by placing the center $c$ at the origin and arranging the leaves evenly on a circle with radius $2r$ around $c$. Clearly, $c$ is adjacent to the leaves. It remains to show that the leaves are pairwise non-adjacent, i.e., they have distance larger than $2r$. Consider two neighboring leaves $u$ and $v$ on the circle and the line segments $cu$ and $cv$. We obtain the isosceles triangle from Figure 7 as in the proof of (ii) with side-length $2r$ and angle $\alpha=\frac{2\pi}{n-1}$. To show that the distance between $u$ and $v$ is greater than $2r$, we have to confirm that $\sin\frac{\alpha}{2}\cdot\sinh 2r>\sinh r$. Using definition of $\sinh$, $r=\log n$, and $\sin x\geq 2x/\pi$ for $x\in[0,\pi/2]$, we obtain
$\sin\frac{\alpha}{2}\cdot\sinh 2r=\sin\frac{\pi}{n-1}\cdot\left(\frac{n^{2}}{2}-\frac{1}{2n^{2}}\right)>n-\frac{1}{n^{3}}>\frac{n}{2}-\frac{1}{2n}=\sinh r,$
which concludes the proof.
@@ 5.1 Covering the separator with cliques
We start by defining the boxes between a hypercycle and its axis $m$ as illustrated in Figure 4. We place points $b_{0},\dots,b_{k}$ along $m$, such that $b_{0}=p$ and they are evenly spaced at distance $\tanh(r)$. Through each point $b_{i}$ we draw a line perpendicular to $m$ and let $a_{i}$ denote its intersection with the upper hypercycle, see Figure 4. We call the region bounded by the line segments $a_{i-1}b_{i-1}$, $b_{i-1}b_{i}$, $b_{i}a_{i}$, and the part of the hypercycle between $a_{i}$ and $a_{i-1}$ a box.
Lemma 5.1.
@@
To calculate the distance from $a_{0}$ to $a_{1}$, we can use that the polygon $a_{0},b_{0},b_{1},a_{1}$ is a Saccheri quadrilateral with base $b_{0}b_{1}$ of length $\tanh r$, legs $b_{0}a_{0}$ and $b_{1}a_{1}$ of length $r$, and summit $a_{0}a_{1}$. The length of the summit is given by the following formula [20, Theorem 10.8]:
@@
since $\sinh r/\cosh r=\tanh r>\tanh 1>\sinh(1/2)$ as $\tanh r$ is monotone increasing. Consequently, $|a_{0}a_{1}|<2r$. Thus, all distances between vertices of a box are at most $2r$, which concludes the proof.
From this lemma, it follows that the vertices inside any box induce a clique. Next, we aim to give an upper bound on the number of boxes we need to cover the separator. For this, let $k$ be the smallest number such that $a_{k}$ lies in the wedge formed by $\ell_{1}$ and $\ell_{2}$. In Figure 4, $k=3$ as $a_{3}$ lies in the gray wedge. Clearly, we can cover the whole separator (blue in Figure 4), with $\mathcal{O}(k)$ boxes on both sides of the line and thus with $\mathcal{O}(k)$ cliques. The following lemma gives a bound on $k$ depending on the opening angle of the wedge.
Lemma 5.3.
@@
Proof 5.4.
Consider the triangle in Figure 4, where $p=b_{0}$ is the apex of the wedge, $d$ is the intersection of the wedge boundary $\ell_{1}$ with the hypercycle, and $c$ is the point on the axis with distance $r$ from $d$. Recall that the distance between $b_{i-1}$ and $b_{i}$ is $\tanh r<1$. Thus, the distance between $b_{0}$ and $b_{k}$ equals the distance $x$ between $p$ and $c$ up to an additive constant. It remains to give an upper bound on $x$.
@@
Using that $\tanh r<1$ and $\tan\varphi>\varphi$, we obtain $\sinh x<\frac{1}{\varphi}$. This yields the claim.